Please help me solve this question.
If the perimeter of a rectangle is 40 centimeters and the legnth is 8 centimeters, determine the width of the rectangle.

Answer by dabanfield(803)

If the perimeter of a rectangle is 40 centimeters and the legnth is 8 centimeters, determine the width of the rectangle.
Let W be the width of the rectangle. Then the perimeter is:
2*width + 2*length = 40
2*W + 2*8 = 40
2*W + 16 = 40
2*W = 40-16
2*W = 24
W = 12